25x = 91 - 16y
16y = 64 - 16x

Answer by checkley71(8403)

25X=91-16Y
-16X=-64+16Y NOW ADD THESE 2 EQUATIONS.
---------------------------------------------
9X=27
X=27/9
X=3 ANSWER.
25*3=91-16Y
75-91=-16Y
-16=-16Y
Y=16/16
Y=1 ANSWER.
PROOF
-16*3=-64+16*1
-48=-64+16
-48=-48